A credit card statement shows an unpaid balance of ₱6,300 on the billing date of September5. Changes and credits to the account are shown below. The finance charge at 2.3% on the average daily balance is entered on the billing date. Find the average daily balance, finance charge and the account balance on the next billing date, which is October 5. Credit Card Transactions: Sept. 10 purchase ₱3,205; Sept. 18 purchase of ₱4,750; Sept. 20 purchase of ₱9,000; and Oct. 1 payment of ₱20,000. Type your answer as: Average Daily Balance = ____________ Finance Charge = __________ Balance owed on October 5 = ____________

1
Expert's answer
2022-02-02T12:05:49-0500

Initially, the balance is equal to 6300, then withdrawals took place, which are indicated with a minus (-3205, -4750,-9000) and then the money was transferred to the card, this amount with a plus: 20,000.

6300+20000-3205-4750-9000=9345


next, we calculate the financial Charge

"9345\times0.023=214.935"


final balance

9345-214.935=9130.07


"Average Daily Balance=\frac{6300+9130.07}{31}=497.74"